# Density of wave fronts

### v2

@@ 6. Polyhedra
We were fascinated to see in the cube case that the wave fronts become more and more complicated and that the wave fronts $W_{t}(P)$ even do lose connectivity. Passing through a corner can cut the wave front because the geodesic flow is no more continuous there. In the universal cover of the cube, once several lattice points have been crossed, the wave front already has had a chance of been hacked into pieces. Computer experiments indicated that the wave front becomes dense even if we start at a corner. There is no Wiederkehr phenomenon also there. The biggest surprise was maybe that the lack of smoothness of the exponential map could leads to situations on polyhedra, where wave fronts become disconnected and complicated already after a relatively short time.
@@
The following result again follows from the corresponding result of the torus. The situation is tougher however. In the square case we dealt with a symmetry group that was the Klein group $G=(a,b,a^{2}=b^{2}=1,ab=ba)$ with 4 elements, which is a commutative group. In the cube case, we deal with the cube group a finite subgroup of $SO(3)$ of order of $24$. Every of the 6 faces of the cube comes in 4 possible rotated orders. If we draw a line $\gamma(t)$ connecting point $P$ with a point $Q$ in a square $X$, then depending on the crossing of the borders we attach a group element to this segment. There is so a rule which assigns to each point $Q$ of the plane an orientation and a point on the surface of the cube. We can model the situation as a 4:1 cover of the open flat manifold in which the 6 corners have been taken away. Imagine that each of the 6 open squares comes in 4 copies. Now these 24 squares can be glued together away from the corners to produce a smooth flat 2 manifold. If we look at the evolution of the wave front $W_{t}\cap Q$ in a face and denote by $W_{t}^{k}$ the by $k\pi/2$ turned front, then we can look at $W_{t}^{0}\cup W_{t}^{1}\cup W_{t}^{2}\cup W_{t}^{3}$ over each square. This union is now the union of 4 wave fronts on the torus. There is now a partition of $S_{t}$ in the universal cover into 4 sets $S_{t}^{k}$ which become increasingly dense each such that $\exp_{t}(S_{t}^{0})$ is the actual wave front $W_{t}(P)$ on the cube. By symmetry if one of the cases $\exp_{t}(S_{t}^{k})$ becomes dense, then all 4 do. Since the union does become dense by the theorem on the torus, also $W_{t}(P)$ does. We have shown:
Corollary 2.
On a cube, all wave fronts become dense.
@@ 7. Existing Literature
The study of wave front propagation in Riemannian geometry intersects with other areas of mathematics like geometric number theory, partial differential equations and dynamical system theory. The case of a torus illustrates the connection with the geometry of numbers or Ehrhart theory, theories that shed light on the distribution and density of lattice points in a solid if the solid is scaled. As an example, we mention Huxley and Nowak’s work on primitive lattice points in convex planar domains [23] that provides a detailed analysis of the distribution of these points within a compact convex subset of $\mathbb{R}^{2}$. There are intricate relationships between the geometry of the domain and the associated error terms. This study offers foundational insights that can be extended to understand how wave fronts might distribute over more complex surfaces. In the geometry of numbers [30], one explores the relationship between geometric structures and lattice point distributions within the context of dynamical systems, particularly focusing on how curvature and boundary shape influence the number of lattice points. Waxman and Yesha’s recently looked at lattice points within thin sectors [39]. How does the number of integer lattice points within shrinking sectors of a circle grow? Such problems are situated at the intersection of analytic number theory and geometric analysis. The Diophantine properties of the slope of the sector’s central line affects the distribution of lattice points. Such studies could help to understand wave front behavior in more constrained geometric settings such as rectangular billiards or on the surface of a cube, where the geometric structure imposes additional complexity.
@@
The problem of wave fronts is also related to the illumination problem: what is the set of points $Q$ such that the wave front $W_{t}(P)$ never reaches $Q$? This question was first asked the 1950ies. Roger Penrose in 1958, constructed the Penrose un-illuminable room. The first polygonal unilluminable room was found in 1995 [35]. In a general polygon, this is not an empty set in general, but it is finite in general [40]. For recent work on translation surfaces see [31]. Note however that having an un-illuminable room does not always prevent that the wave front can become dense. (Added: January 26th: The Penrose example already shows that there are regions for which entire parts of the room can not be illuminated. In that case, also the wave front can not be dense. Thanks to Sergei Tabachnikov to point this out to us.) On the other hand, having a room that can be illuminated does not prove that fronts become dense. The duck in the center of a circular pond illustrates this.
@@
An obvious relation of wave front dynamics is with partial differential equations like wave dynamics. If we have a Riemannian manifold we can look at the wave dynamics on it. If a solution to this partial differential equation starts localized at a point, then the wave front is related to the wave front given by the geodesics. This is related to the Huygens principle. Wave dynamics is also closely linked to quantum mechanics because if $L$ is the Laplacian on the manifold and $L=D^{2}$, then the wave equation $u_{tt}=-Lu$ is by d’Alembert equivalent to two Schrödinger type equations $u_{t}=-iDu$ and $u_{t}=+iDu$ for the square root $D$ of the Laplacian. This motivates semi-classical analysis [14]. There are connections between high energy eigenfunctions of the Laplacian and the geodesic or billiard dynamics [24]. We are not aware of any relation between the wave front spread and the eigenfunctions of the Laplacian.
@@
We have already mentioned that in general, we expect wave fronts to become dense. What about specific situations in higher dimensions like in the case of billiards in a regular polytop? In the case of a polytop that tesselates Euclidean space, we can prove the density by projection. How can one prove for a regular polytop that $W_{t}(P)$ becomes dense? A good place to start would be a solid regular dodecahedron. The proofs given here do not apply. Still, we feel that there should be a simple argument proving that wave fronts are dense in any convex polytop of dimension 2 or larger.

### v4

@@ 6. Polyhedra
We were fascinated to see that in the cube case, the wave fronts became more and more complicated and that the wave fronts $W_{t}(P)$ even lost connectivity. Passing through a corner can cut the wave front because the geodesic flow was no longer continuous there. In the universal cover of the cube, once several lattice points were crossed, the wave front already had a chance of been hacked into pieces. Computer experiments indicated that the wave front becomes dense even if we start at a corner. The biggest surprise was maybe that the lack of smoothness of the exponential map could leads to situations on polyhedra, where wave fronts become disconnected and complicated already after a relatively short time.
Corollary 2.
On the surface $M$ of a cube, all wave fronts become dense.
@@
Fix $P\in M$. It is allowed to be one of the corners of the cube $M$. Let $Q$ be a second point on the cube and assume $\epsilon>0$ is given. We will show that for every $t>144/\epsilon^{2}$, the wave front $W_{t}(P)$ intersects with the $\epsilon$ neighborhood of $Q$. Each $P$ defines a set of measure zero on $M$ such that if $Q$ is not there, then any of the countably many possible geodesic connecting $P$ with $Q$ is defined for all times and does not pass through any of the corners of $M$.
6.4.
The net $N$ of the cube $M$ is defined as a union of 6 translated squares $k,k+1)\times[l,l+1)$ in $\mathbb{R}^{2}$. We can draw it as a union of squares $A,B,C,D,E,F$ in what we call the large torus $0,4)\times[0,4)$. Without loss of generality, the origin of the wave $P$ can be chosen to be in $A$. As there is a folding correspondence between $N$ and $M$, there is a point $Q$ on $N$ that belongs to $Q$ in $M$.
6.5.
Theorem 1 showed that for $3/\sqrt{t}<\epsilon/2$, the wave front $W_{t}(P)$ on the small torus $\mathbb{T}^{2}/\mathbb{Z}^{2}$ covers $\mathbb{T}^{2}$ $(\epsilon/2)$-dense. Scaling implies that for $6/\sqrt{t/4}<\epsilon$, every $(\epsilon/2)$ - neighborhood of $W_{t}(P)$ covers the torus $\mathbb{R}^{2}/(4\mathbb{Z})^{2}$ of side length $4$. This means that for every $t>36/(\epsilon/2)^{2}$, the set $B_{\epsilon/2}(Q)\cap W_{t}(P)$ is not empty. The wave front $W_{t}(P)$ on the large torus is now $\epsilon/2$ dense. It is therefore $(\epsilon/2)$-dense also on each of the squares $\{A,B,C,D,E,F\}$.
$6.$
[…]
6.7.
[…]
Figure 8. To the left, we see the net NN inside the large torus $0,4)\times[0,4)$. The net NN is $M$ that is cut and flattened onto $\mathbb{R}^{2}$. To the right, the color on an eligible line through $P$ encoding the face number of $M$.
@@ 7. Existing Literature
The study of wave front propagation in Riemannian geometry intersects with other areas of mathematics like geometric number theory, partial differential equations and dynamical system theory. The case of a torus also connects with the geometry of numbers or Ehrhart theory. These theories that shed light on the distribution and density of lattice points if the solid is scaled. As an example, we mention Huxley and Nowak’s work on primitive lattice points in convex planar domains [24] that provides a detailed analysis of the distribution of these points within a compact convex subset of $\mathbb{R}^{2}$. There are intricate relationships between the geometry of the domain and the associated error terms. This study offers foundational insights that can be extended to understand how wave fronts might distribute over more complex surfaces. In the geometry of numbers [31], one explores the relationship between geometric structures and lattice point distributions within the context of dynamical systems, particularly focusing on how curvature and boundary shape influence the number of lattice points. Waxman and Yesha’s recently looked at lattice points within thin sectors [40]. How does the number of integer lattice points within shrinking sectors of a circle grow? Such problems are situated at the intersection of analytic number theory and geometric analysis. The Diophantine properties of the slope of the sector’s central line affect the distribution of lattice points. Such studies could help to understand wave front behavior in more constrained geometric settings such as rectangular billiards or on the surface of a cube, where the geometric structure imposes additional complexity.
@@
The problem of wave fronts is also related to the illumination problem: what is the set of points $Q$ such that the wave front $W_{t}(P)$ never reaches $Q$? This question was first asked the 1950s. In 1958, Roger Penrose constructed the Penrose un-illuminable room. In 1995, the first polygonal unilluminable room was found [36]. For a general polygon, this is not an empty set in general, but it is maximally a finite set of points in general [41]. For recent work on translation surfaces see [32]. (Added: January 26th: The Penrose example already shows that there are regions for which entire parts of the room cannot be illuminated, implying the wave front can not be dense. Thanks to Sergei Tabachnikov to point this out to us.) On the other hand, having a room that can be illuminated does not prove that fronts become dense. The duck in the center of a circular pond illustrates this.
@@
An obvious relation of wave front dynamics is with partial differential equations, like wave dynamics. If we have a Riemannian manifold we can look at the wave dynamics on it. If a solution to this partial differential equation starts localized at a point, then the wave front is related to the wave front given by the geodesics. This is related to the Huygens principle. Wave dynamics is also closely linked to quantum mechanics because if $L$ is the Laplacian on the manifold and $L=D^{2}$, then the wave equation $u_{tt}=-Lu$ is by d’Alembert equivalent to two Schrödinger type equations $u_{t}=-iDu$ and $u_{t}=+iDu$ for the square root $D$ of the Laplacian. This motivates semi-classical analysis [13]. There are connections between high energy eigenfunctions of the Laplacian and the geodesic or billiard dynamics [25]. We are not aware of any relation between the wave front spread and the eigenfunctions of the Laplacian.
@@
We have already mentioned that in general, we expect wave fronts to become dense. What about specific situations in higher dimensions, like in the case of billiards in a regular polytop? In the case of a polytop that tesselates Euclidean space, we can prove the density by projection. How can one prove for a regular polytop that $W_{t}(P)$ becomes dense? A good place to start would be a solid regular dodecahedron. The proofs given here do not apply. Still, we feel that there should be a simple argument proving that wave fronts are dense in any convex polytop of dimension 2 or larger.
@@
8. More experiments
8.1.
A referee observed that we have given little numerical evidence for the conjectures. In order to illustrate the situation, we add two pictures. The first is a test case in billiards. It is related to the dodecahedron case. The second is an example of a two-dimensional Riemannian manifold, where we chose a non-flat metric. They are examples of many experiments we have done over the years, both with billiards as well as with geodesic flows.
8.2.
Let us look at the situation of a regular pentagon, where the unfolding produces a Veech surface has genus 2 [16]. Periodic trajectories have slopes in the golden field $\mathbb{Q}(\sqrt{5})$. All orbits are either periodic or uniquely ergodic. We feel that this example could be a case, where the wave front question is accessible. As in any polygon, wave front length is $|W_{t}(P)|=2\pi t$.
Figure 10. We see wave fronts starting at the origin in a regular pentagon of side length $1$. The radii of the fronts are $r=0.66,0.93,1,1.16,3.33,6.66,20$. At $r=1$, when $W_{t}(P)$ reaches a vertex for the first time, this cuts the wave front. For these pictures, we evolved 100’000 billiard trajectories.
8.3.
Lets look now at the situation with a non-flat smooth metric
$g=\left[\begin{array}[]{cc}a^{2}\cos^{2}(u)+1&-a\cos(u)\cos(v)\\ -a\cos(u)\cos(v)&\cos^{2}(v)+1\\ \end{array}\right]$
on the torus $M=\mathbb{T}^{2}=\mathbb{R}^{2}/(2\pi\mathbb{Z})^{2}$. The curvature is $a\sin(u)\sin(v)/(1+a^{2}\cos^{2}(u)+\cos^{2}(v))^{2}$. The pictures show the case with $a=1.2$. We then integrate numerically the geodesic differential equations $\ddot{x}^{k}=-\sum_{i,j}\Gamma_{ij}^{k}\dot{x}^{i}\dot{x}^{j}$ with initial condition $x(0)=p,x^{\prime}(0)=v$ for 100’000 directions.
Figure 11. Wave fronts on a non-flat torus $(M,g)$, where $g$ is a smooth Riemannian metric. The geodesic differential equations define a flow on the unit tangent bundle of $N=M\times\mathbb{T}^{1}$. This system preserves the volume and is in general expected to show sensitive dependence on initial conditions on some part of $N$. In such a case, we would expect the wave front length to grow exponentially.
@@ 8. More experiments
Declarations
8.5.
The authors did not receive support from any organization for the submitted work. No funding was received to assist with the preparation of this manuscript. No funding was received for conducting this study. No funds, grants, or other support was received.
8.6.
All authors whose names appear on the submission made substantial contributions to the conception or design of the work.